OpenStudy (anonymous):

Imagine you put all positive charges from a cubic decimeter of iron on the moon, and the negative charges remain on the earth. How big would the electrical force between the moon and the earth be then?

OpenStudy (anonymous):

I know im gonna use coloumbs law and have to find how many protons and elektrons a cubic decimeter of iron contains.. But how to do it.. thats h ard

OpenStudy (michele_laino):

we have to compute how many atoms of iron there are into 1 dm^3 of iron

OpenStudy (michele_laino):

Now, what is the density of iron?

OpenStudy (anonymous):

7874 kg/m^3

OpenStudy (michele_laino):

ok! Better is 7.874 g/cm^3

OpenStudy (michele_laino):

so waht is the mass contained into 1 dm^3 of iron? keep in mind that: 1 dm^3= 1000 cm^3

OpenStudy (michele_laino):

oops... so what*

OpenStudy (anonymous):

protons and elektrons?

OpenStudy (michele_laino):

no, the mass of iron into 1 dm^3: mass = density * volume

OpenStudy (michele_laino):

mass= 7.874* 1000=...?

OpenStudy (michele_laino):

since: 1 dm^3 = 1000 cm^3

OpenStudy (michele_laino):

mass = 7,874 grams is it ok?

OpenStudy (anonymous):

oh yeah 7874 kg?

OpenStudy (michele_laino):

no, 7874 grams, since the density is 7.874 grams/cm^3

OpenStudy (anonymous):

oh sorry yes ofcourse

OpenStudy (michele_laino):

ok! Now we know that the molecular mass of iron is 55.85. Is it ok?

OpenStudy (michele_laino):

please take a look to your periodic table of elements

OpenStudy (anonymous):

yes thats correct!

OpenStudy (michele_laino):

ok! So we can say that into 1 dm^3 of iron, there are: \[\frac{{7874}}{{55.85}} = 140.98\] moles of iron

OpenStudy (michele_laino):

since 1 mole of iron is 55.85 grams of iron. Is it ok?

OpenStudy (anonymous):

yes!!

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

Now, 1 mole of iron contains 6.022*10^23 atoms of iron. so there are \[140.98 \times 6.023 \times {10^{23}} = 849.12 \times {10^{23}} = 8.5 \times {10^{25}}\] atoms of iron into 1 dm^3 of iron

OpenStudy (michele_laino):

\[\Large 140.98 \times 6.023 \times {10^{23}} = 849.12 \times {10^{23}} = 8.5 \times {10^{25}}\] atoms of iron

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

Now, we have to keep in mind that the atomic number of iron is 26

OpenStudy (michele_laino):

so, the number of protons contained into 1 dm^3 of iron is: \[\Large 8.5 \times {10^{25}} \times 26 = 221 \times {10^{25}} = 2.21 \times {10^{27}}\]

OpenStudy (anonymous):

im with you

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

Now, how many coulombs are \[\Large 2.21 \times {10^{27}}\] protons

OpenStudy (anonymous):

2.21x10^27*1.6x10^-27?

OpenStudy (michele_laino):

I think: \[\Large 2.21 \times {10^{27}} \times 1.6 \times {10^{ - 19}} = ...?\]

OpenStudy (anonymous):

yes ^-19.....

OpenStudy (anonymous):

3.536C?

OpenStudy (michele_laino):

ok!: \[\Large 2.21 \times {10^{27}} \times 1.6 \times {10^{ - 19}} = 3.536 \times {10^8}Coulombs\]

OpenStudy (anonymous):

yea, and that is q_1? and q_1=q_2?

OpenStudy (anonymous):

\[F=k*\frac{ q^2 }{ r^2 }\]

OpenStudy (michele_laino):

Now, we have \[\large 3.536 \times {10^8}Coulombs\] on the moon, and: \[\large 3.536 \times {10^8}Coulombs\] on the earth

OpenStudy (michele_laino):

what is the mean distance moon-earth?

OpenStudy (anonymous):

3,844x10^5m

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

the magnitude F of the requested force, is: \[\Large F = K\frac{{{Q_1}\;{Q_2}}}{{{d^2}}} = 9 \times {10^9} \times \frac{{{{\left( {3.536 \times {{10}^8}} \right)}^2}}}{{{{\left( {3.84 \times {{10}^8}} \right)}^2}}} = ...?\]

OpenStudy (anonymous):

7631406250 N

OpenStudy (michele_laino):

better is: \[\Large F = 7.63 \times {10^9}Newtons\]

OpenStudy (michele_laino):

please note that: \[{Q_1}\; = - {Q_2}\] so the requested force is an attractive force

OpenStudy (anonymous):

oh okey, thanks again! michele, your a saviour!!!

OpenStudy (anonymous):

the

OpenStudy (michele_laino):

thanks! :)

OpenStudy (anonymous):

there were alot of chemistry in this one? in the beginng?

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thats why it was so hard!!

OpenStudy (anonymous):

thanks alot once again! now Ive to go to sleep cya

OpenStudy (michele_laino):

thanks again! :)